Systems Staff Meeting Notes 4/27/98

Systems Staff Meeting Notes
April 27, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Bobbie Parker, Suzy Shaw, Leota Whitfield

Announcements
  1. Destry Fudge will be out until June 15. We probably won’t be seeing him back here, although he has yet to turn in a letter of resignation. His job position cannot be filled until he resigns.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Herd of Cows finally received - 57 of them. There is one stray Cow - destined for Library Office - that has not arrived yet. Seventeen are work stations for staff, bought on auxilliary money. The rest are destined for the public. Bill received a distribution list from Martha on Friday, will try to start moving them today.
Liaisons
  1. Dial-up session was last Thursday - some liaisons are reporting some minor problems - but, overall it was successful.
  2. The demo did work - John Ashcraft & his students scrambled to get it to boot up by 2PM.
  3. Liaisons will be changing staff over to local copies of Office ’97. We have just received 250 individual licenses.
Networking
  1. NT3 - WINS configuration - the print queues in Facilities were gone, now they’re back. It is unknown how the server got out of synch, but it is currently talking to LibPreserve NT just fine.
  2. The option pack to ready for 5.5 was installed on NT2.
  3. NT3 has been shaky since it was made the second WINS server.
  4. Bill needs to reconfigure LibPreserve.
  5. NT3 will have to be reconfigured - 140M full load - to accommodate the 250 Office’97 licenses. Hopefully there will not be a big push to access these all at once - it will slow down the Net. We will need to track the installs for the counts of the licenses. We do have a fax that covers us having secondary licenses on some machines.
  6. Progress on net connection documentation. With the changes in Systems, we discovered that the transceivers feed improves overall access. All of Systems is now fed by a single Boca hub, with the exception of the baby Intel hub in Sharon’s office.
  7. Ghost problem in LAC - added a transceiver to the hub and the problem was fixed. NERDC was interested in this development.
  8. Net & phone wiring is complete in Systems. Phones should be in this week. Physical Plant guy is currently working to configure Special Collections, then he will do Systems. These phones will have a number of different features. There is no message unit per se - but there should be a message light on each phone to indicate that they do, indeed, have a message.
  9. After the phones are completely installed, we will need to move Lee out to her cubicle in Room 10.
  10. Within the next week or so we will make a change in the Education loop - we are just waiting to hear from NERDC. A page on the NERDC website shows that the port usage of the 100M feed is not even at 10%, so the congestion cannot be due to usage. IP capacity will improve.
  11. Preservation - Pre-Bid Conference will be this held this week.
  12. MSL is progressing - the wall should go up this week. However, the furniture order just went out Friday - and it will be 60 - 120 days before it arrives. The Info Swamp II will not be up & running until the furniture arrives. We may have to distribute the Cows we have in the holding area destined for MSL and reorder.
  13. Need to reach a decision on the dumb terminals. The plan was to replace them with IBM network work stations - however, they are still in Beta - and now the price of PC’s has dropped enough to be competitive.
  14. Centurion experiment going well in MSL.
  15. Looks like we’ll go with Windows’98 before NT on the public work stations. Win’98 is due out in June. Microsoft regulatory decisions are still pending.
  16. In Systems we have one Control Unit doing nothing but supporting one printer - the 3270 upstairs. Sharon uses it for patron-updates & Debra uses it for tape work. There are remote print service capabilities available on NT. Some other solution will need to be found, as we want to get rid of the Control Units.
  17. Control Unit in room 112 West can probably be removed. It was supporting one terminal way in the back (Gus Harrer’s) - need to consult with Gary about getting it out.
Software Projects Status
  1. Pre-invoicing will be done on Thursday, April 30 and invoicing on Friday, May 1.
  2. Lending will be off the mainframe by fiscal year end. Borrowing side will still remain on the mainframe.
  3. Sharon is still working on Personnel.
  4. Brian will demo the Patron Tracker for Special Collections this week.
  5. Baldwin - all records have been converted - 10,000 partial records. With WinOni, Baldwin utilization is definitely up.
  6. No change on LIS or Mellon.
  7. Carolyn Henderson decided to go with spreadsheets for the MIS Information circulation.
  8. TSMG - updated list - backdoor LIS functions for tracking in Resource Services.
  9. Suzy encourages us to test the Outlook schedule calendar. It sends out invitations, reserves rooms and gives everyone an individual calendar.
DP Coordinator
  1. Debra has hired 2 new students for the summer term: Stephanie Stults & Becky Leidy.
  2. Jason Krasnove will be returning.
  3. Scott Olipra’s situation is still iffy - he may have to go home to work for his father.
  4. Student work hours will start to diminish by Wednesday because of finals.
  5. Students will begin to return the week of May 11.
Calendar
  1. Sharon out today.
  2. Bill & Suzy out May 14 - May 15.
  3. Debra out May 15.
  4. Memorial Day observed May 25.
  5. Bobbie out June 8 - 19.
  6. Debra out half day June 16, all day June 17.
  7. Bill & Winston out June 15-24.
  8. Independence Day observed July 3.
  9. Debra out July 7 - 10.
  10. Bobbie out July 13 - 17
  11. Brian out July 16 & 17.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu